OpenStudy (cccmpt2):

I can't figure out how to do this problem: One of the factors of 3x^2+x is:

OpenStudy (mathstudent55):

There are two terms: 3x^2 and x The facors of the two terms are: 3x^2 = 3 * x * x x = x Which factor do both terms have in common?

OpenStudy (anonymous):

tahe the common term x(3x+1)

OpenStudy (cccmpt2):

i'm so lost....i don't know how to do the problem at all, i literally have nothing to go from...

OpenStudy (mathstudent55):

What do 3 * x * x and x have in common?

OpenStudy (cccmpt2):

ok, i see how the x is in the first term and x is the second term, so they are equal. so all i have to do is find what is the same factor of the terms?

OpenStudy (mathstudent55):

So the x ic common to both.

OpenStudy (mathstudent55):

That's one of the factors. The question is answered.

OpenStudy (mathstudent55):

Since 3x^2 + x = x(3x + 1), x is one of the factors and 3x + 1 is the other factor.

OpenStudy (cccmpt2):

so in the problem x^2-9, x *x would be a factor and x-9 would be the other, so they would have x in common too?

OpenStudy (mathstudent55):

x^2 - 9 is a special type of factoring called the difference of two squares. a^2 - b^2 = (a+ b)(a - b)

OpenStudy (mathstudent55):

There is no x in common because the second term, -9, has no x.

OpenStudy (cccmpt2):

so i should've factored it as x*x and just -9 which have nothing in common ..?..

OpenStudy (mathstudent55):

In your case, a = x, and b = 9 x^2 - 9 = x^2 - 3^2. As you can see it's the difference of two squares. x^2 is the square of x, and 9 is the square of 3, so x^2 - 9 = (x + 3)(x - 3)

OpenStudy (cccmpt2):

oh wow. i need to study more because this does't make enough sense to me. i really appreciate your help, though!

OpenStudy (cccmpt2):

i just don't understand how (x+3) became part of the answer

OpenStudy (mathstudent55):

When you use FOIL to multiply out (x +3)(x - 3), you do get x^2 - 9.

OpenStudy (cccmpt2):

i can do the FOIL work to check the answer, just can figure out how the (x+3) came to be

OpenStudy (mathstudent55):

When you multiply two binomials that follow this pattern, (a + b)(a - b), the middle term of the soultion cancles out and you get just a^2 - b^2.